Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

both Peter and Paul win

Expert replies
by sanju09 » Fri Jul 04, 2014 10:07 pm
The probability that Peter would win a grill show is twice of the probability that Paul would win the same grill show. If the probability that both Peter and Paul win the grill show is 9/32, what is the probability that only one of the two would win the grill show?
A. ¾
B. 9/16
C. 3/8
D. ¼
E. 5/32


Made Up!
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Sat Jul 05, 2014 3:34 am
sanju09 wrote:The probability that Peter would win a grill show is twice of the probability that Paul would win the same grill show. If the probability that both Peter and Paul win the grill show is 9/32, what is the probability that only one of the two would win the grill show?
A. ¾
B. 9/16
C. 3/8
D. ¼
E. 5/32
Let P(Paul wins) = x.
Since Peter's probability is twice Paul's, P(Peter wins) = 2x.
Since the probability that both win is 9/32, we get:
x * 2x = 9/32.
2x² = 9/32
x² = 9/64
x = 3/8.

Thus:
P(Paul wins) = x = 3/8, implying that P(Paul loses) = 5/8.
P(Peter wins) = 2x = 2 * (3/8) = 3/4, implying that P(Peter loses) = 1/4.

Case 1: P(Paul wins but Peter doesn't) = 3/8 * 1/4 = 3/32.
Case 2: P(Peter wins but Paul doesn't) = 3/4 * 5/8 = 15/32.
Since exactly one person wins in either Case 1 or Case 2, we ADD the fractions:
3/32 + 15/32 = 18/32 = 9/16.

The correct answer is B.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Matt@VeritasPrep » Sun Jul 06, 2014 6:21 pm
You might want to stipulate that the two men's chances of winning are independent, as the GMAC has been dabbling with dependent probability of late, at least going by the Question Pack they just released. (Probability in general seems to be a bigger focus of the test, though the new OG will tell us more.)

Anyway, I thought of this problem as:

(Prob ONLY One Wins) = 1 - (Both Win) - (Neither Wins)
= 1 - (x*2x) - (1-x)(1-2x)
= 3x - 4x²
= x * (3 - 4x)

We're given that 2x² = 9/32, or x = 3/8, so we have

(3/8) * (3 - 3/2), or (3/8) * (3/2), or 9/16.
Join the discussion

by sanju09 » Mon Jul 14, 2014 11:27 pm
Great replies!

If P(Paul wins) = x then P(Peter wins) = 2x; such that (x)(2x) = 9/32 or x = 3/8 (probability that Paul loses is 5/8) and 2x = ¾ (probability that Peter loses is ¼).

Now, the probability that only one of the two would win the grill show means

Peter wins and Paul loses OR Paul wins and Peter loses

= (¾) (5/8) + (3/8)( ¼)

= (15/32) + (3/32)

[spoiler]= 18/32

= 9/16.

Chose (B).
[/spoiler]
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion